OpenStudy (chris215):

The position equation for a particle is s(t)=sqrt(t^3+1) where s is measured in feet and t is measured in seconds. Find the acceleration of the particle at 2 seconds.

OpenStudy (michele_laino):

here we have to find the second derivative of \(s(t)\)

OpenStudy (michele_laino):

namely: \[\huge s''\left( 2 \right)\]

OpenStudy (michele_laino):

for example the first derivative of \(s(t)\), is: \[\Large s'\left( t \right) = \frac{3}{2}\frac{{{t^2}}}{{\sqrt {{t^3} + 1} }}\] now, please compute the second derivative

OpenStudy (chris215):

I got (3 t (t^3+4))/(4 (t^3+1)^(3/2))

OpenStudy (chris215):

My answer is 2/3 ft/sec^2